A number sequence is as follows: 4, 8, 12, 16, 20, 24, .... If 4 is the 1st number in the sequence what is the sum of the first 300 numbers in the sequence?
Answer by fractalier(6550)

Couple of ways to do that...here's one way...
The sum of the first 300 integers: 1,2,3,4,5,... is found by
S = n(n+1)/2 = 300*301/2 = 45150
but since our sequence is four times that, that is, ours goes by 4,8,12,16,...,
we will merely multiply our answer by four and get
S = 4*45150 = 180600